OpenStudy (anonymous):

a gril looks at a boat fromw a light house at a 25degrees angle the boat is beside some rocks 600 feet from the botem of the light house what is her elevation 1287, 280, 254, 544

OpenStudy (anonymous):

can any one help me

OpenStudy (campbell_st):

draw a digram with the necessary information  select a trig ratio sin = opposite/hypotenuse cos= adjacent/hypotenuse tan = opposite/adjacent tan(25) = x/600 so x = 600* tan(25) I'll let you evaluate it